# Superconducting pairing correlations on a trapped-ion quantum computer

### v1

@@ V Conclusion
We note that finite-size corrections to the energy density scaling as $L^{-1}$ for open boundaries and $L^{-2}$ for periodic boundaries is a well-known and generic fact in one-dimensional critical systems, see for example [81]. In the Octagon encoding, the number of qubits $Q$ is related to the linear system size as $L=\sqrt{2Q/5}$. Therefore, to obtain the same amount of finite-size effects in the two settings, one needs
@@
assuming that the ground state can prepared exactly in each case. We note that, for the free fermion case, the fermionic fast Fourier transform can prepare the ground state reasonably efficiently even without periodic boundary conditions being natively available in the hardware (though reconfigurability does confer a benefit even in this case [82]). The quantitative estimate of the advantage of (mixed) periodic qubits shown in Fig. S28c is thus to be seen as a representative example for the simulation of critical states.
@@
For the Hubbard model at half-filling, the quantum Monte Carlo method can provide a numerically exact theory predictions for various physical observables for finite temperatures and the ground state. Such predictions are often used as a metrology tool to determine the temperature of the state prepared in analog quantum simulation [83, 10]. Here, we perform the determinant quantum Monte Carlo calculations on a $6\times 6$ lattice with an interaction strength of $U/t=8$ and all four different boundary conditions, i.e., PBC-PBC, PBC-APBC, APBC-PBC, and APBC-APBC, using the QUEST package [84, 85, 86]. For the PBC-PBC configuration, we employ the default run script “test”, which exploits the C4 rotational symmetry of the lattice. For all other boundary condition setups, simulations are carried out using the “ggeom” run script, which supports a variety of geometric and boundary configurations. Each simulation consists of 20,000 warmup sweeps followed by 80,000 measurement sweeps with a step size $\tau=0.01$. The mixed boundary condition results reported are the average over the simulation results of the four boundary conditions.
@@
This work was made possible by a large group of people, and the authors would like to thank the entire Quantinuum team for their many contributions. We are grateful for helpful discussions with Immanuel Bloch, Harry Buhrman, David Amaro and Cristina Cirstoiu. E.G. acknowledges support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project Bench-QC (DIK0425/01). Some of the matrix-product state simulations were run in compute resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0032628. All experiments were run on Quantinuum’s Helios Quantum computer, powered by Honeywell.
@@
E.G., S-H.L, K.H., R.N., M.I. and H.D. conceived the experimental ideas, implemented and tested the circuits and did the data analysis, with E.G. leading work on the half-filled, S-H.L. on the doped and K.H. on the bilayer model. E.G. developed the error mitigation scheme. A.R., C.F., C.F., J.P.G., A.H., A.H., A.I., C.J.K., N.K., M.M., A.R.M., A.J.P., A.P.R., B.N. and J.G.B. built the experimental apparatus and took the data. J.A. developed circuit debugging tools. D.T.S. developed efficient toric code preparation circuits. R.H., P.A-M., M.F-F., and A.C.P contributed to the classical simulations and to the writing of the manuscript. H.D. drafted the initial manuscript to which all authors contributed.
@@
The data supporting this study is available at Zenodo repository with DOI 10.5281/zenodo.17483197 [87].
@@
H.D. and A.P.R. are shareholder of Quantinuum. All other authors declare no competing interests.

### v2

@@ V Conclusion
We note that finite-size corrections to the energy density scaling as $L^{-1}$ for open boundaries and $L^{-2}$ for periodic boundaries is a well-known and generic fact in one-dimensional critical systems, see for example [pirvu2012matrix]. In the Octagon encoding, the number of qubits $Q$ is related to the linear system size as $L=\sqrt{2Q/5}$. Therefore, to obtain the same amount of finite-size effects in the two settings, one needs
@@
assuming that the ground state can prepared exactly in each case. We note that, for the free fermion case, the fermionic fast Fourier transform can prepare the ground state reasonably efficiently even without periodic boundary conditions being natively available in the hardware (though reconfigurability does confer a benefit even in this case [maskara2025fastsimulationfermionsreconfigurable]). The quantitative estimate of the advantage of (mixed) periodic qubits shown in Fig. S28c is thus to be seen as a representative example for the simulation of critical states.
@@
For the Hubbard model at half-filling, the quantum Monte Carlo method can provide a numerically exact theory predictions for various physical observables for finite temperatures and the ground state. Such predictions are often used as a metrology tool to determine the temperature of the state prepared in analog quantum simulation [mazurenko_cold-atom_2017, xu_neutral-atom_2025]. Here, we perform the determinant quantum Monte Carlo calculations on a $6\times 6$ lattice with an interaction strength of $U/t=8$ and all four different boundary conditions, i.e., PBC-PBC, PBC-APBC, APBC-PBC, and APBC-APBC, using the QUEST package [varney2009quantum, lee2010quest, chang2015recent]. For the PBC-PBC configuration, we employ the default run script “test”, which exploits the C4 rotational symmetry of the lattice. For all other boundary condition setups, simulations are carried out using the “ggeom” run script, which supports a variety of geometric and boundary configurations. Each simulation consists of 20,000 warmup sweeps followed by 80,000 measurement sweeps with a step size $\tau=0.01$. The mixed boundary condition results reported are the average over the simulation results of the four boundary conditions.
@@
This work was made possible by a large group of people, and the authors would like to thank the entire Quantinuum team for their many contributions. We are grateful for helpful discussions with Immanuel Bloch, Harry Buhrman, David Amaro and Cristina Cirstoiu. E.G. acknowledges support by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project Bench-QC (DIK0425/01). At various points during the development of this work, Phasecraft’s compact fermionic encoding, US patent no. 12020119 was used under license. Some of the matrix-product state simulations were run in compute resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility using NERSC award DDR-ERCAP0032628. All experiments were run on Quantinuum’s Helios Quantum computer, powered by Honeywell.
@@
E.G., S-H.L, K.H., R.N., M.I. and H.D. conceived the experimental ideas, implemented and tested the circuits and did the data analysis, with E.G. leading work on the half-filled, S-H.L. on the doped and K.H. on the bilayer model. H.D. defined the high-level experimental protocols for the three settings, developed the injection technique for state preparation and optimized circuit design for the three settings. E.G. developed the error mitigation scheme. R.N. and H.D. worked out the fermionic encodings. A.R., C.F., C.F., J.P.G., A.H., A.H., A.I., C.J.K., N.K., I.S.M., M.M., A.R.M., A.J.P., A.P.R., B.N. and J.G.B. built the experimental apparatus and took the data. J.A. developed circuit debugging tools. D.T.S. developed efficient toric code preparation circuits. R.H., P.A-M., M.F-F., and A.C.P contributed to the classical simulations and to the writing of the manuscript. H.D. drafted the initial manuscript to which all authors contributed.
@@
The data supporting this study is available at Zenodo repository with DOI 10.5281/zenodo.17483197 [zenodo].
@@
H.D. is a shareholder of Quantinuum. All other authors declare no competing interests.